# Synthesizing Programmatic Reinforcement Learning Policies with Large Language Model Guided Search

### v1

@@ Abstract
Programmatic reinforcement learning (PRL) has been explored for representing policies through programs as a means to achieve interpretability and generalization. Despite promising outcomes, current state-of-the-art PRL methods are hindered by sample inefficiency, necessitating tens of millions of program-environment interactions. To tackle this challenge, we introduce a novel LLM-guided search framework (LLM-GS). Our key insight is to leverage the programming expertise and common sense reasoning of LLMs to enhance the efficiency of assumption-free, random-guessing search methods. We address the challenge of LLMs’ inability to generate precise and grammatically correct programs in domain-specific languages (DSLs) by proposing a Pythonic-DSL strategy – an LLM is instructed to initially generate Python codes and then convert them into DSL programs. To further optimize the LLM-generated programs, we develop a search algorithm named Scheduled Hill Climbing, designed to efficiently explore the programmatic search space to consistently improve the programs. Experimental results in the Karel domain demonstrate the superior effectiveness and efficiency of our LLM-GS framework. Extensive ablation studies further verify the critical role of our Pythonic-DSL strategy and Scheduled Hill Climbing algorithm.
@@ 1 Introduction
Deep reinforcement learning (DRL) has achieved great success from beating the world champion in Go [51] to powering the frontier natural language assistants [42, 3], and demonstrated great potential in robotics [28], autonomous vehicles [62], and recommendation systems [11]. However, approximating a policy using a deep neural network makes the decision-making process a black box and, therefore, less interpretable and trustable to human users [25]. Subsequently, carefully designed tools and costly human intervention are often required to deploy trustworthy DRL systems [18]. Moreover, DRL frequently encounters substantial performance declines when applied to previously unseen scenarios [31, 16], revealing another aspect of its challenges in generalization.
@@
Recent programmatic reinforcement learning (PRL) methods have explored representing an RL policy using a program [2, 58, 5, 52]. Instead of learning a state-to-action mapping as in DRL, PRL synthesizes programs written in Domain-Specific Languages (DSLs) as human-readable policies that can be parsed and executed, making human inspection possible as an extra safety guard. Moreover, such structured program policies are shown to be able to capture high-level task-solving ideas, allowing for generalizing to a wide range of task variants [57]. Despite the encouraging results, the state-of-the-art PRL algorithms are notoriously inefficient and require tens of millions of program execution in environments [57, 37, 10]. Under the hood, they are search or RL algorithms without any assumption to the targeting problems. On the one hand, this allows generalization to all kinds of problems; on the other hand, the search time grows exponentially with increasing DSL complexity, making it intolerable for any practical use case.
@@
Our key insight is that there is likely only a limited set of problems corresponding to specific program policies that are of human interest – hence, we can utilize reasonable assumptions to prune the space. Recently, Large Language Models (LLMs) have been demonstrated to possess internet-scale knowledge that can be retrieved by a natural language interface [61, 65, 56]. If we view the text on the internet as a “text projection” of human civilization, an LLM as a knowledge base should be able to provide hints to “prune” the program search paths that are out of the “human interest scope.” With this intuition, we conjecture LLMs can be utilized to bootstrap the sample efficiency of search-based PRL algorithms, pushing PRL one step closer to practical adoption.
@@
To this end, we aim to develop a PRL framework that utilizes LLMs to produce task-solving program policies while minimizing the number of program executions in environments. Directly instructing LLMs to synthesize DSL programs for solving PRL tasks faces three fundamental challenges: (1) LLMs may lack the domain knowledge of the PRL tasks, e.g. environment dynamics, what an agent can obverse or how it can act, (2) the training data of modern LLMs are mostly natural language texts and general-purpose programming languages, e.g. Python and C++, which can be quite different from the DSLs used in PRL, and (3) there is no apparent mechanism for optimizing LLMs to produce programs that maximize rewards since the best-performing LLMs are privately owned.
@@
We compare our proposed LLM-GS framework in the Karel domain to various existing PRL methods [57, 37, 10]. The experimental results demonstrate that LLM-GS is significantly more effective and efficient than the baselines. Extensive ablation studies show that (1) our proposed Pythonic-DSL strategy leads to a higher ratio of executable (i.e. grammatically correct) programs and a higher average return compared to directly generating DSL programs, (2) our proposed search method Scheduled HC achieves the best efficiency among existing search algorithms, (3) initializing the search population of Scheduled HC using LLM-generated programs is significantly more efficient than randomly sampled programs.
@@ 2 Related work
Programmatic reinforcement learning (PRL). PRL represents RL policies using more structured and potentially more interpretable and generalizable representations, such as decision tree [5], state machine [27, 32, 36], symbolic expression [58, 59, 6, 33], Logic programming language [29], and program written in domain-specific language (DSL) [2, 52, 54, 66, 46, 39, 38, 57, 37, 10]. Trivedi et al. [57], Liu et al. [37] devised PRL tasks in the Karel domain [44] and proposed learning embedding spaces of programs using variational autoencoders [30] and then optimizing program embeddings using cross-entropy method or RL. Carvalho et al. [10] achieve the state-of-the-art results in the Karel domain [44] by searching in a programmatic space of AST structure. Despite the encouraging results and the generality of these methods, they are notoriously inefficient, requiring tens of millions of program executions to obtain task-solving programs. In this work, to devise an efficient PRL framework, we make the first attempt to leverage the knowledge and the reasoning ability of LLMs to generate a set of initial programs to bootstrap search algorithms; moreover, we propose Scheduled HC, a search algorithm that best utilizes the LLM-generated programs.
@@
Large language models for code generation. With a remarkable ability to understand and generate natural languages and codes, LLMs have been widely adopted for code generation [7, 1, 40, 64, 48] via neural program synthesis methods [35, 53, 17, 47, 63, 34, 43, 4]. These works target general-purpose programming languages for software development, such as Python and C++, with abundant data on the internet; in contrast, we aim to solve PRL tasks that require writing programs restricted by given domain-specific languages. Prior works exploring using LLMs to synthesize DSL programs via providing LLMs with DSL grammars and few-shot examples, hindsight relabeling, and prioritized experience replay [60, 21, 9]. While these works focus on string and array transformation, and abstract reasoning [15], we leverage LLMs to synthesize DSL program policies to be executed in an RL environment and maximize return.
@@
Search-based program synthesis. Various search algorithms have been developed for program-by-example (PBE) [22, 20, 45, 23, 43, 4], whose goal is to find programs that satisfy given examples, e.g. input/output string pairs. Recent works have explored utilizing search algorithms in DSL to learn libraries [19, 21]. In the regime of programmatic RL, Carvalho et al. [10] recently applied the Hill Climbing algorithm to the Karel benchmark, achieving state-of-the-art performance. Instead of randomly sampling programs via search algorithms, we present a framework that integrates knowledge from an LLM with a search algorithm to significantly improve the sample efficiency.
@@
Figure 1: The Karel DSL grammar. It describes the Karel domain-specific language’s actions, perceptions, and control flows. The domain-specific language is obtained from Liu et al. [37].
@@ 3 Preliminary
The Karel domain. We first review the Karel domain, the de facto test bed for programmatic reinforcement learning research [8, 12, 50, 24, 13]. The Karel domain-specific language illustrated in Figure 2 is a robot programming language to control the Karel agent in a 2D grid world. The agent’s actions include moving as well as interacting with the environment by picking up and putting down objects (markers). The perceptions check for obstacles and markers, which allows observing the environment. Lastly, control flows, e.g. REPEAT, WHILE, IF, and IFELSE, enable describing complex decision-making logics.
We illustrate a Karel task, DoorKey, in Figure 2. The agent needs to explore the left room using the DSL actions such as move and turnLeft as well as the perceptions such as frontIsClear and markersPresent to find the key (marker). After the key is picked up via pickMarker, the wall to the right room will unblock. The agent then needs to place another marker on the goal marker, which appears in the right room, to receive the full reward for the task. Appendix C presents all Karel tasks.
The state-of-the-art performance in Karel is achieved by iteratively searching for improved programs according to their episodic return using search algorithms [10]. Next, we introduce the concept of search space, which defines how search algorithms find neighbor programs from a set of current candidates for improved performance, and then dive into the search algorithm details.
Programmatic space vs. latent space. In the programmatic space, a program $P$ can be represented as an abstract syntax tree (AST), where each leaf node in the AST represents a program token [10]. To obtain its neighborhood program, a node is sampled from its AST, and then replaced by a subtree randomly generated using the DSL’s production rules and sampling strategies. Another choice is to search in a learned latent space [57, 37], constructed by training a variational autoencoder (VAE) on randomly generated DSL programs. To sample a neighbor program, a small noise is added to the latent embedding of the current program candidate and the VAE decoder can generate a program from the noise-corrupted embedding. Appendix B contains more details of search spaces.
@@
Cross-entropy method (CEM). CEM [49] first generates $k$ neighborhood programs from its latent space and evaluates their episodic return on the Karel environment [57]. Next, if the average return over the top-$E$ neighborhood programs is higher than the best-seen average return, it continues the process on the mean latent embedding of the top-$E$ neighborhood programs. When the average return no longer increases, the program with the highest episodic return is returned.
@@ 4 Large language model-guided search for efficient programmatic reinforcement learning
Table 1: Activate LLM’s task awareness. To bridge the domain gap between the target task and the LLM’s knowledge, we craft a carefully curated prompt with relevant knowledge of solving the task without explicitly dictating the specific programming approach or implementation. Take the task DoorKey here as an example: given task name, map description, initial position, and task goal, the LLM is encouraged to solve the specified task with its programming skills and common sense. More details can be found in Section E.2.
@@
The hill climbing algorithm (HC) is the state-of-the-art programmatic reinforcement learning (PRL) approach in Karel, at the cost of tens of millions of program interactions [10]. This prevents its application to real-world decision-making problems, where program-environment interaction at this scale is inapplicable. Hence, we aim to maintain the high episodic return while reducing programs executed in the environments, i.e., improves the sample efficiency of PRL. To this end, our key insight is to utilize a large language model (LLM), hypothesizing that its abundant world knowledge, including programming skills and common sense, may bootstrap the inherently assumption-less, random-guessing HC. However, to implement this idea, many challenges rooted in the environment’s domain gap (Section 4.1), DSL’s language barrier (Section 4.2), and the inability to optimize closed-source LLMs (Section 4.3) must be addressed.
@@ 4.1 Domain and task-aware prompting
An inexperienced LLM user might directly ask an LLM: “Write a program to solve the Karel task DoorKey,” since this is how average users work with LLM-based software development assistants. Obviously, repeating this until the LLM spits out a correct program will not be more efficient than existing search algorithms. The first challenge we encountered was that the LLM lacks background knowledge of the PRL task at hand. After all, there is no guarantee that the LLM has seen or learned the text or code content related to the specific PRL task. Therefore, it is crucial to describe the task and the environment to the LLM in detail using natural language so that it makes the right assumptions based on its learned common sense. Here, we devise a prompting strategy that converts the PRL task’s domain knowledge into natural language sentences that LLMs can process and reason. Table 1 shows the prompt we carefully wrote for Karel DoorKey. Note how the instructions humanly describe the task but do not leak the actual programming skills to implement the required navigation. Now that the LLM possesses minimal knowledge of the task, we may further instruct it to generate programs that solve the desired task.
@@ 4.2 Generating DSL programs with Pythonic-DSL strategy
Despite the issue, we still hope to leverage the best of LLM’s programming skills to synthesize DSL programs. To this end, we derive a solution from the following assumptions: (1) The LLM is proficient in a general-purpose programming language, such as Python, (2) The DSL can be represented by a possibly restricted version of the general-purpose programming language, and (3) The LLM understands the restriction in natural language, if any. Our Pythonic-DSL strategy instructs the LLM to generate Python programs instead, given the Karel rules and constraints we wrote in English, and then later convert it into the Karel DSL. As shown in Table 2, DSL details, including the action, perception functions, and language constraints, are provided. [^1] Additionally, task-agnostic environment knowledge is also given. See Appendix F for the full prompt. Thus, the LLM can generate programs based on its innate general programming skills and its understanding of the DSL description. One caveat is that the LLM still occasionally generates Python programs that cause Python-to-DSL parsing failures. An empirical useful mitigation is instructing the LLM to generate the DSL program converted from its own output Python program as a backup. At this point, we have developed the mechanism to instruct an LLM to generate DSL programs given the PRL task and the DSL grammar in natural language.
Table 2: The Pythonic-DSL instruction. The Karel DSL is specified via a constrained version of Python with pre-defined functions including Pythonic-DSL perceptions that allow the agent to observe, and Pythonic-DSL actions that enable the agent to interact. This Pythonic-DSL description is fed into the LLM via the system prompt so that it can generate Python code that can later be converted into Karel DSL. We will show that this approach outperforms the direct generation of DSL or Python programs in Table 3. The full prompt is presented in Section E.1.

### v2

@@ Abstract
Programmatic reinforcement learning (PRL) has been explored for representing policies through programs as a means to achieve interpretability and generalization. Despite promising outcomes, current state-of-the-art PRL methods are hindered by sample inefficiency, necessitating tens of millions of program-environment interactions. To tackle this challenge, we introduce a novel LLM-guided search framework (LLM-GS). Our key insight is to leverage the programming expertise and common sense reasoning of LLMs to enhance the efficiency of assumption-free, random-guessing search methods. We address the challenge of LLMs’ inability to generate precise and grammatically correct programs in domain-specific languages (DSLs) by proposing a Pythonic-DSL strategy – an LLM is instructed to initially generate Python codes and then convert them into DSL programs. To further optimize the LLM-generated programs, we develop a search algorithm named Scheduled Hill Climbing, designed to efficiently explore the programmatic search space to improve the programs consistently. Experimental results in the Karel domain demonstrate our LLM-GS framework’s superior effectiveness and efficiency. Extensive ablation studies further verify the critical role of our Pythonic-DSL strategy and Scheduled Hill Climbing algorithm. Moreover, we conduct experiments with two novel tasks, showing that LLM-GS enables users without programming skills and knowledge of the domain or DSL to describe the tasks in natural language to obtain performant programs.
@@ 1 Introduction
Deep reinforcement learning (DRL) has achieved great success from beating the world champion in Go (Silver et al., 2016) to powering the frontier natural language assistants (Ouyang et al., 2022; Bai et al., 2022), and demonstrated great potential in robotics (Jain et al., 2024), autonomous vehicles (Wang et al., 2023b), and recommendation systems (Chen et al., 2023). However, approximating a policy using a deep neural network makes the decision-making process a black box and, therefore, less interpretable and trustable to human users (Heuillet et al., 2021). Subsequently, carefully designed tools and costly human intervention are often required to deploy trustworthy DRL systems (Doshi-Velez and Kim, 2017). Moreover, DRL frequently encounters substantial performance declines when applied to previously unseen scenarios (Kirk et al., 2023; Cobbe et al., 2018), revealing another aspect of its challenges in generalization.
@@
Recent programmatic reinforcement learning (PRL) methods have explored representing an RL policy using a program (Andre and Russell, 2001; Verma et al., 2018; Bastani et al., 2018; Silver et al., 2020). Instead of learning a state-to-action mapping as in DRL, PRL synthesizes programs written in Domain-Specific Languages (DSLs) as human-readable policies that can be parsed and executed, making human inspection possible as an extra safety guard. Moreover, such structured program policies are shown to be able to capture high-level task-solving ideas, allowing for generalizing to a wide range of task variants (Trivedi et al., 2021). Despite the encouraging results, the state-of-the-art PRL algorithms are notoriously inefficient and require tens of millions of program execution in environments (Trivedi et al., 2021; Liu et al., 2023; Carvalho et al., 2024). Under the hood, they are search or RL algorithms without any assumption to the targeting problems. On the one hand, this allows generalization to all kinds of problems; on the other hand, the search time grows exponentially with increasing DSL complexity, making it intolerable for any practical use case.
@@
Our key insight is that there is likely only a limited set of problems corresponding to specific program policies that are of human interest – hence, we can utilize reasonable assumptions to prune the search space. Recently, large language models (LLMs) have been demonstrated to possess internet-scale knowledge that can be retrieved by a natural language interface (Wang et al., 2024; Zheng et al., 2024; Taylor et al., 2022). If we view the text on the internet as a “text projection” of human civilization, an LLM as a knowledge base should be able to provide hints to “prune” the program search paths that are out of the “human interest scope”. With this intuition, we conjecture LLMs can be utilized to bootstrap the sample efficiency of search-based PRL algorithms, pushing PRL one step closer to practical adoption.
@@
To this end, we aim to develop a PRL framework that utilizes LLMs to produce task-solving program policies while minimizing the number of program executions in environments. Directly instructing LLMs to synthesize DSL programs for solving PRL tasks faces three fundamental challenges: (1) LLMs may lack the domain knowledge of the PRL tasks, e.g., environment dynamics, what an agent can obverse or how it can act, (2) the training data of modern LLMs are mostly natural language texts and general-purpose programming languages, e.g., Python and C++, which can be quite different from the DSLs used in PRL, and (3) there is no apparent mechanism for directly and iteratively optimizing LLMs to produce programs that maximize rewards since the best-performing LLMs are privately owned, e.g., GPT-4 (Achiam et al., 2023).
@@
We compare our proposed LLM-GS framework in the Karel domain to various existing PRL methods (Trivedi et al., 2021; Liu et al., 2023; Carvalho et al., 2024). The experimental results demonstrate that LLM-GS is significantly more effective and efficient than the existing methods. Extensive ablation studies show that (1) our proposed Pythonic-DSL strategy leads to a higher ratio of executable (i.e. grammatically correct) programs and a higher average return compared to directly generating DSL programs, (2) our proposed search method, Scheduled HC, achieves the best efficiency among existing search algorithms, (3) initializing the search population of Scheduled HC using LLM-generated programs is significantly more efficient than randomly sampled programs. To evaluate whether LLM-GS is useful to users without knowledge of the Karel domain and DSL, we additionally design two novel tasks and only provide LLM-GS with task descriptions while fixing the domain and DSL prompts. The experiment results show that LLM-GS still achieves significantly improved sample efficiency, highlighting the extensibility of our proposed framework.
@@ 2 Related work
Programmatic reinforcement learning (PRL). PRL represents RL policies using more structured and potentially more interpretable and generalizable representations, such as decision tree (Bastani et al., 2018), state machine (Inala et al., 2020; Koul et al., 2019; Lin et al., 2024), symbolic expression (Verma et al., 2018; 2019; Bhupatiraju et al., 2018; Landajuela et al., 2021), Logic programming language (Jiang and Luo, 2019), and program written in domain-specific language (DSL) (Andre and Russell, 2001; Silver et al., 2020; Sun et al., 2020; Zhu et al., 2019; Qiu and Zhu, 2022; Moraes and Lelis, 2024; Mariño et al., 2021; Trivedi et al., 2021; Liu et al., 2023; Carvalho et al., 2024). Trivedi et al. (2021); Liu et al. (2023) devised PRL tasks in the Karel domain (Pattis, 1981) and proposed learning embedding spaces of programs using variational autoencoders (Kingma and Welling, 2014) and then optimizing program embeddings using cross-entropy method or RL. Carvalho et al. (2024) achieve the state-of-the-art results in the Karel domain (Pattis, 1981) by searching in a programmatic space of AST structure. Despite the encouraging results and the generality of these methods, they are notoriously inefficient, requiring tens of millions of program executions to obtain task-solving programs. In this work, to devise an efficient PRL framework, we make the first attempt to leverage the knowledge and the reasoning ability of LLMs to generate a set of initial programs to bootstrap search algorithms; moreover, we propose Scheduled HC, a search algorithm that best utilizes the LLM-generated programs.
@@
Large language models for code generation. With a remarkable ability to understand and generate natural languages and codes, LLMs have been widely adopted for code generation (Brown et al., 2020; Achiam et al., 2023; Nijkamp et al., 2022; Xu et al., 2022; Roziere et al., 2023) via neural program synthesis methods (Lin et al., 2018; Sun et al., 2018; Desai et al., 2016; Raza et al., 2015; Wang et al., 2017; Le et al., 2022; Parisotto et al., 2017; Balog et al., 2017). These works target general-purpose programming languages for software development, such as Python and C++, with abundant data on the internet; in contrast, we aim to solve PRL tasks that require writing programs restricted by given domain-specific languages. Prior works exploring using LLMs to synthesize DSL programs via providing LLMs with DSL grammars and few-shot examples, hindsight relabeling, and prioritized experience replay (Wang et al., 2023a; Grand et al., 2024; Butt et al., 2024). While these works focus on string and array transformation as well as abstract reasoning (Chollet, 2019), we leverage LLMs to synthesize DSL program policies to be executed in an RL environment and maximize the return.
@@
Search-based program synthesis. Various search algorithms have been developed for program-by-example (PBE) (Gulwani, 2011; Feser et al., 2015; Polozov and Gulwani, 2015; Gulwani et al., 2017; Parisotto et al., 2017; Balog et al., 2017), whose goal is to find programs that satisfy given examples, e.g. input/output string pairs. Recent works have explored utilizing search algorithms in DSL to learn libraries (Ellis et al., 2021; Grand et al., 2024). In the regime of programmatic RL, Carvalho et al. (2024) recently applied the Hill Climbing algorithm to the Karel benchmark, achieving state-of-the-art performance. Instead of randomly sampling programs via search algorithms, we present a framework that integrates knowledge from an LLM with a search algorithm to significantly improve the sample efficiency.
An extended discussion on related work can be found in Appendix A.
@@
Figure 1: The Karel DSL grammar. It describes the Karel domain-specific language’s actions, perceptions, and control flows. The domain-specific language is obtained from Liu et al. (2023).
@@ 3 Preliminary
The Karel domain. We first review the Karel domain, the de facto test bed for programmatic reinforcement learning research (Bunel et al., 2018; Chen et al., 2019; Shin et al., 2018; Gupta et al., 2020; Chen et al., 2021). The Karel domain-specific language illustrated in Figure 2 is a robot programming language to control the Karel agent in a 2D grid world. The agent’s actions include moving as well as interacting with the environment by picking up and putting down objects (markers). The perceptions check for obstacles and markers, which allows observing the environment. Lastly, control flows, e.g., REPEAT, WHILE, IF, and IFELSE, enable describing complex decision-making logics. More details of the DSL can be found in Appendix B.
We illustrate a Karel task, DoorKey, in Figure 2. The agent needs to explore the left room using the DSL actions such as move and turnLeft as well as the perceptions such as frontIsClear and markersPresent to find the key (marker). After picking up the key via pickMarker, the wall to the right room will unblock. The agent then needs to place a marker on the goal marker located in the right room to receive a full task reward. Appendix C presents all the Karel tasks in detail.
The state-of-the-art performance in Karel is achieved by iteratively searching for improved programs according to their episodic return using search algorithms (Carvalho et al., 2024). Next, we introduce the concept of search space, which defines how search algorithms find neighbor programs from a set of current candidates for improved performance, and then dive into the search algorithm details.
Programmatic space vs. latent space. In the programmatic space, a program $P$ can be represented as an abstract syntax tree (AST), where each leaf node in the AST represents a program token (Carvalho et al., 2024). To obtain its neighborhood program, a node is sampled from its AST, and then replaced by a subtree randomly generated using the DSL’s production rules and sampling strategies. Another choice is to search in a learned latent space (Trivedi et al., 2021; Liu et al., 2023), constructed by training a variational autoencoder on randomly generated DSL programs. To sample a neighbor program, a small noise is added to the latent embedding of the current program candidate, and the decoder can generate a program from the noise-corrupted embedding.
@@
Cross-entropy method (CEM). CEM (Rubinstein and Kroese, 2004) first generates $k$ neighborhood programs from its latent space and evaluates their episodic return on the Karel environment (Trivedi et al., 2021). Next, if the average return over the top-$E$ neighborhood programs is higher than the best-seen average return, it continues the process on the mean latent embedding of the top-$E$ neighborhood programs. When the average return no longer increases, the program with the highest episodic return is returned.
@@ 4 Large language model-guided search for efficient programmatic reinforcement learning
Table 1: Activate LLM’s domain and task awareness. To bridge the domain gap between the target task and the LLM’s knowledge, we curate a scaffolding prompt to alleviate the knowledge gap and activate the task-solving ability without explicitly dictating the specific programming approach or implementation. Take the task DoorKey as an example, the user writes down the task name, goal, and some detailed information like map and initial position. Users can easily follow the categories in the prompt to write the task description. The LLM is encouraged to solve the specified task with its programming skills and common sense. More details of the prompts can be found in Appendix D.
@@
The hill climbing algorithm (HC) is the state-of-the-art programmatic reinforcement learning (PRL) approach in Karel, at the cost of tens of millions of program interactions (Carvalho et al., 2024). This prevents its application to real-world decision-making problems, where program-environment interaction at this scale is inapplicable. Hence, we aim to maintain the high episodic return while reducing programs executed in the environments, i.e., improves the sample efficiency of PRL. To this end, our key insight is to utilize a large language model (LLM), hypothesizing that its abundant world knowledge, including programming skills and common sense, may bootstrap the inherently assumption-less, random-guessing HC. However, to implement this idea, many challenges rooted in the environment’s domain gap (Section 4.1), DSL’s language barrier (Section 4.2), and the inability to optimize closed-source LLMs (Section 4.3) must be addressed. Figure 3 presents an overview of our proposed framework, large language model-guided search (LLM-GS).
@@ 4.1 Domain and task-aware prompting
An inexperienced LLM user might directly ask an LLM: “Write a program to solve the Karel task DoorKey,” since this is how typical users interact with LLM-based software development assistants. Obviously, repeating this until the LLM spits out a correct program is unlikely to be more efficient than existing search algorithms. There are two major challenges that exist: (1) the LLM lacks the dynamic environmental concepts inherent to the Karel domain, and (2) the LLM does not have background knowledge of the specific PRL task DoorKey. Therefore, it is crucial to provide a detailed description of the task and environment in natural language so the LLM can make informed assumptions based on its learned common sense.
To address this, we devise a prompting strategy that converts the task description and the task-agnostic environment knowledge into natural language sentences that LLMs can process and reason. The task-agnostic environment knowledge prompt introduces the basic mechanics of interacting with this environment to the model, setting the stage for more specific programming tasks within the Karel framework, detailed in Section D.1. Table 1 provides an example of a user prompt for the Karel task DoorKey. Now that the LLM has acquired basic knowledge of the Karel domain and the task, we can further instruct it to leverage its programming skills, algorithmic knowledge, and long-term planning abilities to generate programs that solve tasks. Note that given this domain-aware prompt, given any novel task within the same domain, even a user without any programming skills can simply describe the task in natural language to obtain a performant program, as shown in Section 5.4.
@@ 4.2 Generating DSL programs with Pythonic-DSL strategy
Despite the issue, we still hope to leverage the best of LLM’s programming skills, algorithmic knowledge, and long-term planning abilities to synthesize DSL programs. To this end, we derive a solution from the following assumptions: (1) The LLM is proficient in a general-purpose programming language, such as Python, (2) The DSL can be represented by a possibly restricted version of the general-purpose programming language, and (3) The LLM understands the restriction in natural language, if any. Our Pythonic-DSL strategy instructs the LLM to generate Python programs instead, given the Karel rules and constraints we wrote in English, and then later convert it into the Karel DSL. As shown in Table 2, DSL details are provided, including the action, perception, primitives and language constraints. Thus, the LLM can generate programs based on its innate general programming skills and its understanding of the DSL description. One caveat is that the LLM still occasionally generates Python programs that cause Python-to-DSL parsing failures. An empirical useful mitigation is instructing the LLM to generate the DSL program converted from its own output Python program as a backup. At this point, we have developed the mechanism to instruct an LLM to generate DSL programs given the PRL task and the DSL grammar in natural language.
Table 2: The Pythonic-DSL instruction. The Karel DSL is specified via a constrained version of Python with pre-defined functions including Pythonic-DSL perceptions that allow the agent to observe, and Pythonic-DSL actions that enable the agent to interact. This Pythonic-DSL description is fed into the LLM via the system prompt so that it can generate Python code that can later be converted into Karel DSL. We will show that this approach outperforms the direct generation of DSL or Python programs in Table 3. The full prompt is presented in Section D.1.